OpenStudy (anonymous):

9. The average height of female students in a college in the state of NH is 63 inches, while the variance is 53 inches. This statement is true. True False

OpenStudy (anonymous):

std is sqrt of variance = sqrt(53)= about 7. in population wise continuous uniform distribution, std is +- 3 from the mean. minimum height will then be 63 - 3*7= 63-21= 42. makes sense. I would say true.

OpenStudy (anonymous):

assuming that acceptable range is about 1% double tail.